# Catch Me if You Search: When Contextual Web Search Results Affect the Detection of Hallucinations

### v1

@@ 3.5 Search Results in the Static and Dynamic Conditions
We analyzed the search results across static and dynamic conditions to assess whether differences in search result quality contributed to our findings. Each search result was systematically annotated for information such as relevance, correctness, and NewsGuard rating. NewsGuard is a service that evaluates the reliability of web sources based on expert analysis and provides credibility scores to assess the trustworthiness of online content (NewsGuard, ). However, comparing the search results between the static and dynamic conditions revealed no statistically significant differences. We additionally compared the static and dynamic search results across three content categories: genuine content, minor hallucinations, and major hallucinations and observed no significant differences for any content category. Nevertheless, the retrieved results demonstrated strong relevance to the posed questions (static: 92.59%, dynamic: 92.05%) and the majority of the results contained at least one correct answer (static: 90.74%, dynamic: 90.76%), indicating that the search results accurately addressed the questions. Please refer to C for a detailed analysis of the search results in the static and dynamic conditions.
@@ 4.1.1 Effect of Contextual Web Search Results on Truth Discernment
One might ask if participants rated the genuine content as more accurate in the dynamic condition because they were presented with search results better aligned with the LLMs’ responses than those provided in the static condition. To evaluate this possibility, we compared the search results in the static and dynamic conditions for genuine content, minor hallucination, and major hallucination. We found no significant differences across content categories in terms of relevance, correctness, NewsGuard ratings, or any other measured variables. Thus, we can eliminate the possibility that different search results account for different ratings of genuine content between the dynamic and static conditions. Instead, the very act of searching for the answer, rather than mere exposure to search results, seems to have triggered different cognitive processes.
@@ Appendix C Analyses of Search Results in the Static and Dynamic Conditions
After the annotation phase, one annotator used the information in the “Link to the source”columns to compute the NewsGuard reliability ratings, used in prior work to evaluate the reliability of web sources (Aslett et al., 2024). Ratings of 0-39 indicate proceeding with maximum caution, 40-59 indicate proceeding with caution, 60-74 indicate credibility with exceptions, 75-99 indicate generally credible, and 100 indicate high credibility, as outlined in the NewsGuard website (NewsGuard, ). Notably, many sources lacked a defined NewsGuard rating and were rated as “Undefined”, including several seemingly reliable sources, such as the USDA Food Safety and Inspection Service, the Stanford Solar Center, and the World Bank.
Although the search queries for participants in the dynamic condition were pre-entered, they had the flexibility to modify them. However, 97.38% of the search queries remained unchanged. Static and dynamic search results were comparable, with no significant differences, as depicted in Figure 9 (a). We observed similarities in terms of whether they were relevant (static: 92.59%, dynamic: 92.05%), correct (static: 64.81%, dynamic: 65.53%), wrong (static: 9.26%, dynamic: 8.55%), contained a Gen-AI response (static: 53.49%, dynamic: 51.31%), contained at least one correct answer (static: 90.74%, dynamic: 90.76%), or NewsGuard ratings (Figure 9 (b)), measured using chi-square tests with significance levels of 0.05. In addition, 63.54% of dynamic search results duplicated search results from the static condition. We additionally compared the static and dynamic search results for genuine content, minor hallucinations, and major hallucinations and found no significant differences.

### v2

@@ 3.5 Search Results in the Static and Dynamic Conditions
We analyzed the search results across static and dynamic conditions to assess whether differences in search result quality contributed to our findings. Each search result was systematically annotated for information such as relevance and correctness. However, comparing the search results between the static and dynamic conditions revealed no statistically significant differences. We additionally compared the static and dynamic search results across three content categories: genuine content, minor hallucinations, and major hallucinations and observed no significant differences for any content category. Nevertheless, the retrieved results demonstrated strong relevance to the posed questions (static: 92.59%, dynamic: 92.05%) and the majority of the results contained at least one correct answer (static: 90.74%, dynamic: 90.76%), indicating that the search results accurately addressed the questions. Please refer to C for a detailed analysis of the search results in the static and dynamic conditions.
@@ 4.1.1 Effect of Contextual Web Search Results on Truth Discernment
One might ask if participants rated the genuine content as more accurate in the dynamic condition because they were presented with search results better aligned with the LLMs’ responses than those provided in the static condition. To evaluate this possibility, we compared the search results in the static and dynamic conditions for genuine content, minor hallucination, and major hallucination. We found no significant differences across content categories in terms of relevance, correctness, or any other measured variables. Thus, we can eliminate the possibility that different search results account for different ratings of genuine content between the dynamic and static conditions. Instead, the very act of searching for the answer, rather than mere exposure to search results, seems to have triggered different cognitive processes.
@@ Appendix C Analyses of Search Results in the Static and Dynamic Conditions
Although the search queries for participants in the dynamic condition were pre-entered, they had the flexibility to modify them. However, 97.38% of the search queries remained unchanged. Static and dynamic search results were comparable, with no significant differences, as depicted in Figure 9 (a). We observed similarities in terms of whether they were relevant (static: 92.59%, dynamic: 92.05%), correct (static: 64.81%, dynamic: 65.53%), wrong (static: 9.26%, dynamic: 8.55%), contained a Gen-AI response (static: 53.49%, dynamic: 51.31%), or contained at least one correct answer (static: 90.74%, dynamic: 90.76%) (Figure 9 (b)), measured using chi-square tests with significance levels of 0.05. In addition, 63.54% of dynamic search results duplicated search results from the static condition. We additionally compared the static and dynamic search results for genuine content, minor hallucinations, and major hallucinations and found no significant differences.